An airplane can travewl 380 miles into the wind in the same time that it can travel 420 miles with the wind. If the wind speed is 10 miles per hour, find the speed of the airplane without any wind.
Answer by mananth(16949)

plane speed x mph
wind speed 10 mph
..
against wind x-10mph
with wind x+10 mph

Distance against wind 380
Distance with tail wind 420
Time = same
380/(x- 10)= 420/(x+ 10)
380*(x+ 10)= 420*(x- 10)
380x +3800 =420x-4200
420x -380x =4200+3800
40x =8000
/40
x= 200 mph